# Naturalistic Music Decoding from EEG Data via Latent Diffusion Models

### v2

@@ 1 Introduction
Recent works [30, 31] based on Stable Diffusion [12] have enabled the synthesis of high-quality images reconstructed from EEG recordings (for a more informative survey about the growing literature on multi-modal brainwave decoding, the reader can consult [32]). In this article we explore the use of ControlNet [33] as a mechanism for controlling AudioLDM2 [17], a large-scale pre-trained diffusion model on audio, using EEG data. While such a mechanism has been explored in the EEG to image literature with the CMVDM model [34], to our knowledge we are the first work that explores ControlNet for EEG to audio decoding (music in particular), performing such a task (i) without manual pre-processing/filtering and (ii) leveraging a naturalistic music dataset [35].
@@ 2.1 Latent Diffusion Models
In the following we review latent diffusion models [12, 36], on which we base our methodology. Given a VAE [37] mapping an audio signal (e.g., waveform or mel-spectrogram) $\mathbf{x}$ to a latent representation $\mathbf{z}$ via a stochastic encoder encoder $E^{\text{VAE}}$, i.e., $\mathbf{z}\sim E^{\text{VAE}}(\mathbf{x})$, we can define a prior $p_{\theta}$ via a diffusion model $\epsilon_{\theta}$. Based on the DDPM formulation [10], $\epsilon_{\theta}$ is trained to revert a forward Gaussian process, indexed by a time variable $t$ ranging in $[0,T]$:
@@ 2.2 ControlNet
initialized with the weights of a pre-trained $E_{\theta_{E}}$, where $P$ is a projection matrix mapping $\mathbf{y}$ to a tensor having the dimensions of $\mathbf{z}_{t}$ and $c_{\text{in}}$ a zero initialized convolution. While the original ControlNet adapter is additionally conditioned on text (following $E_{\theta_{E}}$), we omit text-conditioning given that we do not use such a modality in this work. The ControlNet adapter outputs, for each layer $i=1,\dots,I$, a feature $C^{i}_{\phi}(\mathbf{z}_{t},\mathbf{y},t)$ which is first convolved with a zero-initialized convolution $c_{i}$, then summed to the features of the original U-Net encoder $E^{i}_{\theta}(\mathbf{z}_{t},t)$, defining:
@@ 3.1 EEG-Conditioned ControlNet
In our method, we employ the ControlNet adapter to condition a (pre-trained) diffusion model on raw EEG information $\mathbf{y}\in\mathbb{R}^{F_{\mathbf{y}}\times S_{\mathbf{y}}}$, where $F_{\mathbf{y}}$ is the number of EEG channels and $S_{\mathbf{y}}$ the number of sampled time-steps. In this work, $E^{\text{VAE}}$ is a 2D convolutional encoder mapping mel-spectrograms $\mathbf{x}\in\mathbb{R}^{F_{\mathbf{x}}\times S_{\mathbf{x}}}$ ($F_{\mathbf{x}}$ is the number of frequencies and $S_{\mathbf{x}}$ the number of bins) to the latent representation $\mathbf{z}\in\mathbb{R}^{F_{\mathbf{z}}\times S_{\mathbf{z}}\times D_{\mathbf{z}}}$, where $F_{\mathbf{z}}$ and $S_{\mathbf{z}}$ are the compressed window dimensions and $D_{\mathbf{z}}$ the latent channels. As shown in Eq. 3, we need to map $\mathbf{y}$ with a projector $P$ to a tensor having the dimensions of $\mathbf{z}$ in order to combine the two sources of information in the ControlNet adapter input. We define the projector $P:\mathbb{R}^{F_{\mathbf{y}}\times S_{\mathbf{y}}}\to\mathbb{R}^{F_{\mathbf{z}}\times S_{\mathbf{z}}\times D_{\mathbf{z}}}$ as a strided 1D conv-net applied over the temporal dimension of $\mathbf{y}$, followed by a reshape operation.
@@ 3.2 Neural Embedding-Based Metrics
  • Pearson Coefficient — Given a generated track $\hat{\mathbf{x}}$ and a ground truth track $\mathbf{x}$, the (sample) Pearson correlation coefficient $r$ is defined as:
@@
Model (Architecture - Subject ) CLAP EnCodec Test set OOD Track Test set OOD Track Test set OOD Track
Conv-2 1.09 320.78 0.26 - -0.0032 - 3.03 -
Pretrained 0.80 163.31 0.38 0.38 0.010 0.0047 1.93 1.91
@@
ControlNet-All 0.42 147.06 0.59 0.39 0.010 0.0074 1.99 1.82
Scratch-2 0.43 158.24 0.55 0.60 0.003 0.0019 2.26 1.95
Scratch-All 0.36 174.65 0.60 0.50 0.007 0.0043 2.21 2.06
@@ 4.1 Dataset
We train our models on NMED-T (Naturalistic Music EEG Dataset - Tempo) [35]. The dataset contains the EEG recordings of 20 adult subjects listening, each, to 10 high-quality real songs (not synthesized from MIDI or monotimbral). The EEG data is recorded with the Electrical Geodesics, Inc. (EGI) GES300 system [46] using 128 channels, and it is sampled at 1 kHz.
For our experiments, we keep the first track (Trigger 21) as a held-out, out-of-distribution test track. The other 9 songs are divided into train (0-80%), validation (80-90%), and test (90-100%) subchunks. The models are validated with the CLAP score during training. Such a choice stems from the difficulty of doing early stopping with the score matching metric in Eq. 2 during validation. Our minimal pre-processing pipeline involves excluding the face channels (retaining 124 channels), which we center using the mean over the first 1000 time steps, and we set 20 standard deviations for clamping.
@@ 4.2 Implementation Details
We implement the proposed ControlNet on top of the AudioLDM2 diffusion repository. Specifically, we use the audioldm2-music checkpoint, publicly available on HugginFace [^1] https://huggingface.co/docs/diffusers/api/pipelines/audioldm2. For the ControlNet projection, we use (256, 512, 1024, 2048) channels with strides (5, 2, 2, 2). All models are trained with Adam [47] optimizer using a learning rate of $10^{-4}$. During training and inference, we ignore the textual conditioning mechanism of AudioLDM2 using as an input the constant string “Pop music”.
@@ 5 Experiments
For our experiments, we train two types of models. First, we train the ControlNet with the frozen diffusion model, both only using data of subject 2 (ControlNet-2) and using all subject’s data (ControlNet-All). At the same time, we want to test how the model performs when training both the ControlNet adapter and the diffusion U-Net from scratch, leading to the Scratch-2 and Scracth-All models. We compare the models both with Conv-2, a simple convolutional network trained on subject 2 data (following the structure of the projector $P$), and the “Pretrained” diffusion model, generating audio unconditionally (i.e., no EEG conditioning is occurring). When evaluating “All” models, we always use subject 2 test data.
We showcase qualitative examples in Figure 2, and on our site [^2] https://emilianpostolache.com/brainwave, and list quantitative results in Table 1. The OOD Track represents the held-out track (see Section 4).
As we can see, in all metrics, the proposed method outperforms the regressor convolutional baseline (which was trained also on the OOD track). From the Table, we can gather additional information. We observe that the MSE (EnCodec) metric is not optimal for evaluation: the pre-trained model, which does not have access to EEG data, apparently beats our models trained specifically for the task. The situation improves with the Pearson Coefficient (EnCodec) metric. Here, the ControlNet-2 model achieves the best performance, yet Pretrained seems to be on par with the ControlNet-All version, still suggesting difficulties in evaluating the models.
Moving to global metrics such as the CLAP Score and the FAD, we see a higher improvement with respect to the Conv-2 and Pretrained baselines, reaching a 0.60 CLAP Score on ControlNet-2 and Scrach-All. On FAD, as well, our models show improved performance. Nonetheless, these metrics still cannot capture entirely the behaviour of the model given that it assigns a very high CLAP Score on the OOD Track with Scratch-2. Given the scarcity of data (we train only using 9 songs), we can only reconstruct unknown segments of known tracks. Aggregating the different metrics, we can agree that ControlNet-2 seems to perform best.
In Figure 3, we report the confusion matrices related to a classification experiment similar to the one in [6], both using the fused and unfused versions of the CLAP embedder. As the reader can observe, our model resembles more closely a diagonal matrix with respect to Conv-2.
@@ 6 Conclusions
In this paper, we have investigated the use of ControlNet for the task of naturalistic music decoding from EEG brainwaves. While we obtain non-negligible decoding results on unknown segments of known tracks, further research is required to improve generalization to distribution shift, both in terms of larger datasets and improved algorithms.

### v3

@@ 1 Introduction
Recent works [30, 31] based on Stable Diffusion [12] have enabled the synthesis of high-quality images reconstructed from EEG recordings (for a more informative survey about the growing literature on multi-modal brainwave decoding, the reader can consult [32]). In this article, we explore the use of ControlNet [33] as a mechanism for controlling AudioLDM2 [17], a large-scale pre-trained diffusion model on audio, using EEG data. While such a mechanism has been explored in the EEG to image literature with the CMVDM model [34], to our knowledge we are the first work that explores ControlNet for EEG to audio decoding (music in particular), performing such a task (i) without manual pre-processing/filtering and (ii) leveraging a naturalistic music dataset [35].
@@ 2.1 Latent Diffusion Models
In the following, we review latent diffusion models [12, 36], on which we base our methodology. Given a VAE [37] mapping an audio signal (e.g., waveform or mel-spectrogram) $\mathbf{x}$ to a latent representation $\mathbf{z}$ via a stochastic encoder $E^{\text{VAE}}$, i.e., $\mathbf{z}\sim E^{\text{VAE}}(\mathbf{x})$, we can define a prior $p_{\theta}$ via a diffusion model $\epsilon_{\theta}$. Based on the DDPM formulation [10], $\epsilon_{\theta}$ is trained to revert a forward Gaussian process, indexed by a time variable $t$ ranging in $[0,T]$:
@@ 2.2 ControlNet
initialized with the weights of a pre-trained $E_{\theta_{E}}$, where $P$ is a projection network mapping $\mathbf{y}$ to a tensor having the dimensions of $\mathbf{z}_{t}$ and $c_{\text{in}}$ a zero initialized convolution. While the original ControlNet adapter is additionally conditioned on text (following $E_{\theta_{E}}$), we omit text-conditioning given that we do not use such a modality in this work. The ControlNet adapter outputs, for each layer $i=1,\dots,I$, a feature $C^{i}_{\phi}(\mathbf{z}_{t},\mathbf{y},t)$ which is first convolved with a zero-initialized convolution $c_{i}$, then summed to the features of the original U-Net encoder $E^{i}_{\theta}(\mathbf{z}_{t},t)$, defining:
@@ 3.1 EEG-Conditioned ControlNet
In our method, we employ the ControlNet adapter to condition a (pre-trained) diffusion model on raw EEG information $\mathbf{y}\in\mathbb{R}^{F_{\mathbf{y}}\times S_{\mathbf{y}}}$, where $F_{\mathbf{y}}$ is the number of EEG channels and $S_{\mathbf{y}}$ the number of sampled time-steps. In this work, $E^{\text{VAE}}$ is a 2D convolutional encoder mapping mel-spectrograms $\mathbf{x}\in\mathbb{R}^{F_{\mathbf{x}}\times S_{\mathbf{x}}}$ ($F_{\mathbf{x}}$ is the number of frequencies and $S_{\mathbf{x}}$ the number of time bins) to the latent representation $\mathbf{z}\in\mathbb{R}^{F_{\mathbf{z}}\times S_{\mathbf{z}}\times D_{\mathbf{z}}}$, where $F_{\mathbf{z}}$ and $S_{\mathbf{z}}$ are the compressed window dimensions and $D_{\mathbf{z}}$ the latent channels. As shown in Eq. 3, we need to map $\mathbf{y}$ with a projector $P$ to a tensor having the dimensions of $\mathbf{z}$ in order to combine the two sources of information in the ControlNet adapter input. We define the projector $P:\mathbb{R}^{F_{\mathbf{y}}\times S_{\mathbf{y}}}\to\mathbb{R}^{F_{\mathbf{z}}\times S_{\mathbf{z}}\times D_{\mathbf{z}}}$ as a strided 1D conv-net applied over the temporal dimension of $\mathbf{y}$, followed by a reshape operation.
@@ 3.2 Neural Embedding-Based Metrics
  • Pearson Coefficient — Given embeddings of a ground truth track $\mathbf{e}$ and of a generated track $\hat{\mathbf{e}}$, the (sample) Pearson correlation coefficient $r$ is defined as:
@@
To estimate the statistical significance of the metrics, we also compute $p$-values using the $z$-test, sampling 500 misaligned pairs of ground truth and generated audios for the null hypothesis.
@@
Experiment CLAP EnCodec Test set OOD Track Test set OOD Track Test set OOD Track
Conv-2 1.09 320.78 0.26 $(p=0.33)$ 0.26 $(p=0.10)$ 3.19 $(p=0.44)$
Pretrained 0.80 163.31 0.38 $(p=0.91)$ 0.38 $(p=0.16)$ 0.63 $(p=0.77)$ 0.64 $(p=0.13)$ 2.35 $(p=0.86)$ 2.18 $(p=0.27)$
@@
ControlNet-All 0.42 147.06 0.59 $\mathbf{(p=1.64\cdot 10^{-10})}$ 0.39 $(p=0.08)$ 0.65 $(p=0.33)$ 0.65 $(p=0.86)$ 2.15 $(p=0.58)$ 1.94 $(p=0.74)$
Scratch-2 0.43 158.24 0.55 $\mathbf{(p=1.35\cdot 10^{-5})}$ 0.60 $(p=0.70)$ 0.61 $(p=0.40)$ 0.63 $(p=0.50)$ 2.46 $(p=0.63)$ 2.14 $(p=0.49)$
Scratch-All 0.36 174.65 0.60 $\mathbf{(p=1.51\cdot 10^{-9})}$ 0.50 $(p=0.94)$ 0.63 $(p=0.29)$ 0.63 $(p=0.19)$ 2.42 $(p=0.52)$ 2.21 $(p=0.17)$
@@ 4.1 Dataset
We train our models on NMED-T (Naturalistic Music EEG Dataset - Tempo) [35]. The dataset contains the EEG recordings of 20 adult subjects listening, each, to 10 high-quality real songs (not synthesized from MIDI or monotimbral). The EEG data is recorded with the Electrical Geodesics, Inc. (EGI) GES300 system [46] using 128 channels, and it is sampled at 1 kHz. For our experiments, we keep the first track (Trigger 21) as a held-out, out-of-distribution (OOD) test track. The other 9 songs are divided into train (0-80%), validation (80-90%), and test (90-100%) subchunks. The models are validated with the CLAP score during training. Such a choice stems from the difficulty of doing early stopping with the score matching metric in Eq. 2 during validation. Our minimal pre-processing pipeline involves excluding the face channels (retaining 124 channels), which we center using the mean over the first 1000 time steps and setting 20 standard deviations for clamping.
@@ 4.2 Implementation Details
We implement the proposed ControlNet on top of the AudioLDM2 diffusion repository. Specifically, we use the audioldm2-music checkpoint, publicly available on HuggingFace [^1] https://huggingface.co/docs/diffusers/api/pipelines/audioldm2. For the ControlNet projection, we use (256, 512, 1024, 2048) channels with strides (5, 2, 2, 2). All models are trained with Adam [47] optimizer using a learning rate of $10^{-4}$. During training and inference, we ignore the textual conditioning mechanism of AudioLDM2 using as an input the constant string “Pop music”. For all metrics in Table 1 we use the unfused version of CLAP.
@@ 5 Experiments
For our experiments, we train two types of models. First, we train the ControlNet with the frozen diffusion model, using both data from subject 2 (ControlNet-2) and all subjects’ data (ControlNet-All). At the same time, we want to test how the model performs when training both the ControlNet adapter and the diffusion U-Net from scratch, leading to the Scratch-2 and Scracth-All models. We compare the models both with Conv-2, a simple convolutional network trained on subject 2 data (following the structure of the projector $P$), and the “Pretrained” diffusion model, generating audio unconditionally (i.e., no EEG conditioning is occurring). When evaluating “All” models, we always use subject 2 test data.
[…]
On all metrics, the proposed method outperforms the regressor convolutional baseline (which was also trained on the OOD track). We observe that the Pearson Coefficient and the MSE metrics computed with EnCodec embeddings are not optimal for evaluation: the pre-trained model, which does not have access to EEG data, scores very similarly to our models. This is also reflected in the $p$-values, which do not show statistical significance in these cases (assuming a threshold of $0.05$). Moving to global metrics (CLAP Score and FAD), we see a higher improvement with respect to the Conv-2 and Pretrained baselines, reaching a 0.60 CLAP Score on ControlNet-2 and Scrach-All. On FAD, our models also show improved performance. With CLAP Score on the test set, the $p$-values show statistical significance. Notice that while we often obtain higher results on the OOD Track experiments (especially with EnCodec-based metrics), such numbers have high $p$ values, and as such, such metrics do not have high statistical significance.
In Figure 3, we report the confusion matrices related to a classification experiment similar to the one in [6], both using the fused and unfused versions of the CLAP embedder. As the reader can observe, our model resembles a diagonal matrix more closely than Conv-2.
@@ 6 Conclusions
In this paper, we have investigated the use of latent diffusion models for the task of naturalistic music decoding from EEG brainwaves. While we obtain non-negligible decoding results on unknown segments of known tracks, further research is required to improve generalization to distribution shift, both in terms of larger datasets and improved algorithms.